Find the width of a rectangle with a diagonal of 17 cm and a length of 15 cm? Using pythagorean theorem.
Answer by cleomenius(959)

A^2 + B^2 = c^2
15^2 + x^2 = 17^2
225 + x^2 = 289
x^2 = 64
x = 8
Cleomenius.